Calculate each of the following quantities in 2.50 moles H3PO4:

atoms of P
1
Expert's answer
2020-10-14T05:55:19-0400

Given the molicular formula = H3PO4

Hence, no. Of 'P' atom in each molicule of H3PO4 is = 1


1 mole = 6.023×10^23 molicules

2.50 mole =( 2.50× 6.023×10^23 )molicules

= 15.057 × 10^23 molicules


So,

In 1 molicule of H3PO4, 1 atom of P

In (15.057 × 10^23) of H3PO4

( 15.057 × 10^23) atoms of P



Hence, in 2.50 moles of H3PO4 there are 15.057 × 10^23 atoms of P